### v2

# This work has been submitted to the IEEE Access for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible. xx.xxxx/ACCESS.2024.xxxxxxx

Abstract
To adapt to real-world dynamics, intelligent systems need to assimilate new knowledge without catastrophic forgetting, where learning new tasks leads to a degradation in performance on old tasks. To address this, continual learning concept is proposed for enabling autonomous systems to acquire new knowledge and dynamically adapt to changing environments. Specifically, energy-efficient continual learning is needed to ensure the functionality of autonomous systems under tight compute and memory resource budgets (i.e., so-called autonomous embedded systems). Neuromorphic computing, with brain-inspired Spiking Neural Networks (SNNs), offers inherent advantages for enabling low-power/energy continual learning in autonomous embedded systems. In this paper, we comprehensively discuss the foundations and methods for enabling continual learning in neural networks, then analyze the state-of-the-art works considering SNNs. Afterward, comparative analyses of existing methods are conducted while considering crucial design factors, such as network complexity, memory, latency, and power/energy efficiency. We also explore the practical applications that can benefit from SNN-based continual learning and open challenges in real-world scenarios. In this manner, our survey provides valuable insights into the recent advancements of SNN-based continual learning for real-world application use-cases.
Index Terms:
Brain-inspired neuromorphic computing, spiking neural networks, event-based processing, continual learning, autonomous embedded systems.
[…]
=-21pt
I Introduction
Humans continuously learn from their interactions with the environment throughout their lifetime [1], exhibiting remarkable cognitive flexibility of brains to enable simultaneous learning and retention of multiple skills. For instance, a person fluent in English can also master Spanish and French, each with unique grammatical structures and vocabulary, without losing proficiency in their native language. This ability is rooted in the brains’ delicate balance between plasticity and stability [2, 3]. Plasticity allows for the acquisition of new skills, while stability ensures that existing knowledge is preserved. Naturally, we expect Artificial Intelligence (AI) systems [4, 5, 6], wielding Neural Network (NN) algorithms [7], to develop a similar learning capability to effectively operate and adapt in real-world scenarios. While state-of-the-art AI systems excel in single task-based static environments, such as image recognition [8], face identification [9], and speech recognition [10], they often struggle in dynamic environments where data or tasks may change in structure, distribution, or characteristics over time. For instance, when a new task is encountered after a resource-intensive training, NNs are often retrained from scratch despite the high computational costs [11]. Therefore, robustness to multiple tasks and sequential experiences, remains a significant research challenge for AI systems. When faced with incremental learning of different tasks, most NNs underperform due to suffering from rapid performance degradation, a phenomenon known as Catastrophic Forgetting (CF) or interference [12, 13, 14].
In recent years, Continual Learning (CL) [15, 16, 17] emerges as a conceptual solution for addressing CF in AI systems. CL aims to balance the system’s ability to learn new tasks without forgetting the previous knowledge, known as the stability-plasticity dilemma [18, 19] (Fig. 1). It also seeks to achieve generalizability across tasks. In addition to avoiding CF, the primary objectives of CL also include ensuring scalability of NNs, minimizing reliance on old data, realizing controlled forgetting, and enabling rapid adaptation and recovery, which will be further discussed in Section II-B.
Figure 1: The key challenge of performing sequential task learning in NNs is balancing stability and plasticity of the weights. Failing to maintain this balance results in CF, leading to a significant performance decline on earlier tasks. It is illustrated by the sharp drop in accuracy on the previous “cat” task as the network learns a new task, such as recognizing “dog”; adapted from studies in [20].
Numerous CL methods have been proposed to address CF in the conventional Artificial/Deep Neural Network (ANN/DNN) domain, which showed notable performance improvements [21, 22, 23, 24, 25, 26, 27]. However, most methods are resource-intensive, demanding additional memory and computational power. Therefore, DNN-based CL methods often fail to account for storage usage during incremental training, necessitating a significant addition of memory for parameters [28]. These conditions are not suitable for autonomous embedded systems, that use portable-battery and need quick adaptation to new data within limited resources. However, designing resource- and energy-efficient systems with CL capabilities (i.e., CL systems) is still a major challenge.
Recently, the brain-inspired Neuromorphic Computing (NC) paradigm [29, 30, 31, 32, 33] has emerged as a promising field for enabling efficient and low-power information processing and decision-making, i.e., by emulating the architecture and functionality of brains with Spiking Neural Network (SNN) algorithms [34, 35, 36, 37, 38, 39]. The event-driven nature of SNNs supports energy-efficient learning of both static and non-static data streams [34, 35, 36, 37, 38, 39]. Furthermore, SNNs can perform unsupervised learning due to their bio-plausible learning rules (e.g., Spike-Timing-Dependent Plasticity (STDP) [40]), thus enabling continuous adaptation to dynamically changing environments and efficiently learning spatio-temporal data online without labels [41, 42]. In this manner, NC aligns seamlessly with the desired characteristics (desiderata) of CL, called Neuromorphic Continual Learning (NCL).
I-A Surveys on Continual Learning (CL)
To fully understand the foundations and recent advancements in CL, a comprehensive survey is imperative. Most of the existing CL surveys [16, 17, 43, 44, 45, 46, 47, 48, 49, 50, 51, 11, 52] have particularly focused on addressing CF in DNNs. Studies of [16] provided an up-to-date CL survey in DNNs. Other surveys focused on categorizing CL methods [43, 47, 44, 49], studying the challenges of forgetting in DNNs [11, 45, 50], reviewing computational requirements of CL (e.g., regularization) [11], discussing the benefits of forgetting in CL [45], providing experimental comparisons of 11 CL methods [50], discussing Online Continual Learning (OCL) paradigm [51], and discussing the importance of adopting brain-inspired data representations [53]. Although all these surveys provide insightful theoretical foundations and information regarding the recent advancements in CL for DNNs, most of them have not addressed CL with bio-plausible SNNs.
The remarkable CL capability of humans has sparked interest in the emerging of NCL paradigm, i.e., investigation on how SNN architectures and learning rules can be utilized to develop efficient CL systems. However, notable surveys on NCL still focused on exploring biological mechanisms for CL [34], and providing the background theories [54]. Therefore, they have not provided a holistic survey that provides a comprehensive view of NCL, ranging from theories to applications. Toward this, in this paper, we provide an in-depth review on preliminaries of CL and NC, state-of-the-art advancements of CL research in ANNs/DNNs and SNNs, and real-world applications for NCL (Fig. 2). Moreover, we also discuss the open research challenges, hence offering an updated perspective for future research directions in developing energy-efficient and adaptive AI systems. A comparison summary of our survey and the existing ones is presented in Table I.
Figure 2: Overview of our survey: CL in ANNs/DNNs vs. SNNs (i.e., NCL). DNNs rely on global learning, supervised training, and computationally intensive processing on conventional hardware, thus making them less suitable for OCL. In contrast, SNNs leverage event-based processing, local learning rules, and low-power neuromorphic hardware which enable energy-efficient CL, thus making them suitable for OCL.
TABLE I: A qualitative comparison of our survey and the existing ones. Note: “✓”, “+”, and “-” mean full, partial, and no consideration, respectively.
Work Year Research Objective Taxonomy Hardware Applications
Questions CL NCL Platforms
[17] 2019 - - - - -
[44] 2021 - - - +
[50] 2021 - - - -
[46] 2022 - - + - -
[47] 2022 - - + - +
[48] 2022 - - + - -
[49] 2022 - - + - +
[52] 2022 - - + - + -
[51] 2022 - - -
[43] 2023 - - - -
[45] 2023 - - - -
[34] 2023 - - - +
[53] 2023 - + - -
[11] 2024 - - + - -
[16] 2024 - - -
[54] 2024 + - -
Our Survey 2024
I-B Our Novel Contributions
Our survey addresses the key questions in developing CL with neuromorphic computing (SNNs), as the following.
  • What are the desiderata, evaluation scenarios, and performance metrics for CL?
  • What are the existing CL methods in the ML field to overcome CF, and how can they be systematically categorized?
  • What are the limitations of the existing CL methods?
  • Why are the energy-efficient CL algorithms crucial for resource-constrained autonomous embedded systems?
  • What techniques do the state-of-the-art propose to enable energy-efficient CL with SNNs (i.e., NCL)?
  • What are the real-world application use-cases that benefit from NCL?
  • What are the open challenges to be considered for enabling energy-efficient NCL?
By addressing the above key questions, our paper makes the following major novel contributions:
  1. 1.
    We introduce the fundamentals of CL problems by explaining its basic formulation, desiderata, and possible learning scenarios. Furthermore, we highlight the OCL paradigm and its importance for practical applications.
  2. 2.
    We discuss the existing methods for enabling CL, which are pre-dominantly from DNN domain. Afterward, we provide different possible CL settings, and provides metrics for evaluating CL.
  3. 3.
    We discuss the background of NC (SNNs), then provide an up-to-date comprehensive review of state-of-the-art works for NCL.
  4. 4.
    We survey a range of real-world applications that benefit from the energy-efficient NCL. Toward the end, we discuss the open research challenges to provide a perspective on future research directions.
Figure 3: Organization of our survey paper
I-C Paper Structure
The rest of this survey paper is organized as shown in Fig. 3. Section II delves into the CL fundamentals. Section III provides an overview of SNNs and the state-of-the-art NCL methods. In Section IV, a summary of real-world application use-cases of NCL is presented. Then, Section V discusses open research challenges for meeting the CL desiderata with SNNs. Finally, Section VI presents the conclusion.
II Continual Learning
In recent years, DNNs have revolutionized the field of AI [55, 51, 56], due to their effective training technique (i.e., gradient descent-based backpropagation), while assuming the training data are independently and identically distributed (i.i.d) [57, 58]. However, a critical question arises: how can we integrate new data into existing models? One approach is to retrain with new data, while using current parameters as the initial state [58]. However, successive training tends to cause NNs to suffer from CF [59, 60]. Although retraining from scratch effectively tackles CF issues, but it is inefficient and almost impossible in some cases, such as real-time learning scenarios where the model must update and adapt to continuous data streams instantly with limited resources (e.g., developmental learning for autonomous agents [61]).
CL (also known as incremental learning or lifelong learning) addresses CF by accumulating knowledge from a continuous data stream throughout their lifetime [1]. As stability-plasticity dilemma requires trade-off between preserving past knowledge (stability) and adapting to new experiences (plasticity[62], researchers have developed several CL methods to provide the same capability [63, 64, 65]. While many CL methods focused on the offline learning with supervised task-based incremental learning, which operates under the assumption of i.i.d. data and constant task identification, these assumptions often diverge from real-world scenarios. In practical applications, input data streams are typically not i.i.d. and task identity may not be available. It highlights the complexity of CL problems, showing the importance of understanding the CL foundations. Therefore, in this section, we delve into the basic formulation and desiderata of CL, and explore various learning scenarios, methods, settings, and evaluation metrics.
II-A Basic Formulation
Figure 4: (a) A robotic arm being trained to perform a variety of tasks sequentially, and is subsequently able to select from its repertoire of learned skills to apply in different situations. (b) Key desiderata for CL; adapted from studies in [1].
A CL algorithm must adapt to new tasks without full access to previous training data while maintaining performance on respective test sets. Formally, defined by parameters $\theta=\cup_{t=1}^{\mathcal{T}}\theta^{(t)}$, where $\theta^{(t)}=\{e^{(t)},\psi\}$, $e^{(t)}$ is the task-specific parameters and $\psi$ is the task-sharing parameters for a task t. A batch of training samples for a task t is denoted as $\mathcal{D}_{t,b}=$$\mathcal{X}_{t,b}$, $\mathcal{Y}_{t,b}$} where $\mathcal{X}_{t,b}$ represents input data, $\mathcal{Y}_{t,b}$ denotes data labels, and t $\in$ $\mathcal{T}$ = {1,···,$\mathcal{T}$} signifies the task identity, with b $\in$ $\mathcal{B}_{t}$ as the batch index ($\mathcal{T}$ and $\mathcal{B}_{t}$ indicate their respective spaces). Task t is defined by its training samples $\mathcal{D}_{t}$, with distribution $\mathbb{D}_{t}$ :=assign:=:= p($\mathcal{X}_{t},\mathcal{Y}_{t}$), $\mathcal{D}_{t}$ encompasses the entire training set, omitting the batch index and assuming no distribution discrepancy between training and test sets. However, in practical scenarios, data labels $\mathcal{Y}_{t}$ and task identity (Task ID) t may not always be available. CL accommodates varying batch sizes for each task’s training samples (i.e., $\mathcal{D}_{t,b}\}_{b\in\mathcal{B}_{t}}\}_{t\in\mathcal{T}}$) or simultaneous arrival (i.e., $\mathcal{D}_{t}\}_{t\in\mathcal{T}}$) [16].
II-B Desiderata of CL
In this section, we explain the desired characteristics (desiderata) of a CL algorithm, and show an overview of key requirements to achieve CL capability using a robotic arm as an example in Fig. 4.
II-B1 Scalability
A CL algorithm should effectively train on a large or potentially unlimited number of tasks without expanding NNs excessively. The increasing of NN capacity and computational costs have to be minimal or sub-linear, thereby ensuring the scalability of a CL algorithm [53]. To achieve high scalability, the major challenges include scalability of network model size [16, 66] and scalability of regularization terms (e.g., weight regularization [67]).
II-B2 No/Minimal Usage of Old Data
Minimizing or eliminating the reliance on data from previous tasks is crucial for a CL algorithm, because of storage constraints and privacy concerns. Recent researches have made progress to achieve this goal [68, 69, 70, 71, 72, 73]. For instance, previous studies utilized mutual information maximization and compressed gradients to minimize the reliance on old data [68, 69, 70], and employed representative subsets to preserve knowledge [71].
II-B3 Task Agnostic (TA)
A CL algorithm should function independently from pre-defined task boundaries during the training and inference phases, known as Task Agnostic (TA). For instance, a robot in an ever-changing environment should adapt to new tasks, such as picking up different objects or navigating new terrains without explicit information of task identity. It has to learn from environment and experiences, and dynamically adjust to new tasks. State-of-the-art attempted to achieve TA using methods studied in [74, 75, 76, 77, 78, 79, 80, 81].
II-B4 Positive Forward Transfer (FWT)
A CL algorithm should leverage previous knowledge to enhance performance on new tasks, known as Positive Forward Transfer (FWT). Formally, $\text{FWT}_{k}$ evaluates the average influence of all old tasks on the current $k$-th task, and can be stated as:
$\text{FWT}_{k}=\frac{1}{k-1}\sum_{j=2}^{k}(a_{j,j}-\tilde{a}_{j})$ (1)
where, $\tilde{a}_{j}$ is the accuracy of a randomly-initialized model trained with data $\mathcal{D}_{j}$ for the $j$-th task [16]. Recent researches explored different methods to enable FWT, including studies in [82, 83, 84, 85, 86, 87, 88, 53, 89, 90, 91].
II-B5 Positive Backward Transfer (BWT)
A CL algorithm should transfer knowledge from later tasks to past tasks for improving performance on the past tasks, known as Positive Backward Transfer (BWT). Here, zero backward transfer suggests zero forgetting [53]. Formally, $\text{BWT}_{k}$ evaluates the average influence of learning the $k$-th task on all old tasks, and can be stated as:
$\text{BWT}_{k}=\frac{1}{k-1}\sum_{j=1}^{k-1}(a_{k,j}-a_{j,j})$ (2)
where, $a_{k,j}$ is the accuracy of a model at $j$-th task after incremental learning of the $k$-th task [16].
II-B6 Controlled Forgetting
A CL algorithm should forget old yet insignificant knowledge to make room for learning new information, known as controlled forgetting. Although humans do not experience sudden memory loss, a gradual decrease in memory (forgetting) over time is natural, hence controlled forgetting is beneficial for managing memory retention [92]. Related methods include [93, 94, 95] for DNNs and [96, 97, 41, 42] for SNNs.
II-B7 Fast Adaptation and Recovery
A CL algorithm should quickly learn new tasks while minimizing the loss of previous knowledge (fast adaptation), and regain previous performance after encountering new tasks that cause degradation in learned knowledge (recovery). Related methods include regularization [67] and gradient-based meta-learning or fast optimization [98].
II-C Scenarios in CL
CL scenarios can be categorized using a taxonomy presented in Fig. 5, while their formal comparison is provided in Table II. The initially-identified scenarios are Task-Incremental Learning (Task-IL), Domain-Incremental Learning (Domain-IL), and Class-Incremental Learning (Class-IL). While these scenarios cover some desiderata of CL, they fail to meet the requirement of TA [53]; see Fig. 6. They assume that tasks have clear and well-defined boundaries during training. Therefore, recent studies proposed expanding the CL scenarios by introducing Discrete TA (DTA) [74], Continuous TA (CTA) [74], and Open-World TA (OWTA) [53].
Figure 5: Taxonomy of the CL scenarios based on the nature of training and inference setups considering the task identity (Task ID); adapted from studies in [53].
[…]
CL Difference between $\mathcal{D}_{t-1}$ and $\mathcal{D}_{t}$ Task ID Online
Scenario
p$(\mathcal{X}_{t-1})\neq$ p$(\mathcal{X}_{t})$
p$(\mathcal{Y}_{t-1})\neq$ p$(\mathcal{Y}_{t})$
$\{\mathcal{Y}_{t-1}\neq$ $\mathcal{Y}_{t}$}
Train Test Learning
Task-IL Known Known No
Domain-IL $\times$ $\times$ Known Not required Optional
Class-IL $\times$ Known To be inferred Optional
DTA $\times$ $\times$ Not required Not required Yes
OWTA $\times$ To be inferred To be inferred Yes
CTA $\times$ $\times$ $\times$ Unknown Unknown Yes
Figure 6: Split MNIST according to Task-IL, Domain-IL, and Class-ID. (a) Split MNIST is obtained by splitting the original MNIST into five contexts/tasks, each having two classes. (b) Overview of what is expected of the model at test time for each scenario.
II-C1 Task-Incremental Learning (Task-IL)
[…]
II-C2 Domain-Incremental Learning (Domain-IL)
In Domain-IL, a model learns to solve the same problem in different tasks [99, 100, 101]. Tasks have the same data label space $\{\mathcal{Y}_{t-1}=\mathcal{Y}_{t}\}$ but different input distributions $\textit{p}(\mathcal{X}_{t-1})\neq\textit{p}(\mathcal{X}_{t})$. The model does not require Task IDs in the inference phase, as each task has the same possible outputs (e.g., same classes in each task). The use of a single-headed model (i.e., same output head for every task) ensures that the output space remains consistent. An example of Task-IL is incrementally learning to recognize objects midst varying lighting conditions (e.g., indoor and outdoor) [102].
II-C3 Class-Incremental Learning (Class-IL)
In Class-IL, a model must distinguish the global labels (classes) [99, 100, 101]. Here, Task IDs are only provided in the training phase, as they will be inferred in the inference phase. Due to the multi-class property, $\textit{p}(\mathcal{Y}_{t-1})\neq\textit{p}(\mathcal{Y}_{t})$ is a natural consequence, thereby posing more challenges as compared to Task-IL and Domain-IL.
II-C4 Discrete Task Agnostic (DTA)
[…]
II-C5 Open-World Task Agnostic (OWTA)
In OWTA, a model learns from a sequence of distinct tasks, but needs to infer the Task ID during training and inference phases [53, 74]. It requires the CL algorithm to handle an open set of tasks and classes, potentially facing unknown categories that are not encountered during initial training. This makes the OWTA as one of the most challenging scenarios.
II-C6 Continuous Task Agnostic (CTA)
[…]
II-D Online Continual Learning (OCL)
In OCL, a model learns from a continuous data stream, so that each input sample during inference is used as training data for updating the models’ knowledge [103, 104, 46]. As a new sample arrives, it is immediately used for inference, and for updating the model on-the-fly. This capability is highly desired for systems that face highly dynamic environments (e.g., autonomous mobile agents), where new data is encountered in real-time and immediate updates to the model are required. Its challenges include integrating new knowledge and retaining previous information, while considering limited memory and power/energy budgets. In this survey, we emphasize that OCL is the expected capability for autonomous embedded systems.
II-E CL Methods
In this section, we discuss existing methods for addressing CF in various CL scenarios. These methods are categorized into five approaches based on representation, regularization, rehearsal/replay, optimization, and architecture [16], as follows.
II-E1 Representation-based Approach
This approach is characterized by exploiting the strengths of data representations [16]; see Fig. 7. The related methods are as follows.
Figure 7: An illustration of the taxonomy of (a) representation-based approach and (b) its respective methods; adapted from studies in [16].
  1. 1.
    Self-Supervised Learning (SSL): It makes the model learn to generate useful data representations without relying on explicit labels, by creating supervisory signals from the input data to uncover the underlying structure of the data [105, 106]. Related studies are discussed in [107, 108, 109, 110, 111, 112, 113, 114, 115].
  2. 2.
    Pre-training for Downstream CL: It is the initial phase of model training on a large diverse dataset to learn general-purpose representations, which can be fine-tuned for specific downstream tasks, hence having strong knowledge transfer [83, 116, 117, 118]. To maintain generalizability for future tasks, several strategies have been developed, such as Pre-trained Representations [119, 120, 88, 121, 122], Task-adaptive Prompts [16, 123, 124, 125, 126, 127, 128], Saving Prototypes and Enhancing Classifiers [129, 130, 131], and Optimizing an Updatable Backbone [132, 133, 83, 134, 135].
  3. 3.
    Continual Pre-Training (CPT): It includes techniques studied in [136, 137, 138, 139, 140, 141, 142].
II-E2 Regularization-based Approach
This approach is characterized by adding explicit regularization terms to balance the old and new tasks [16], which usually requires storing a frozen copy of the old model for reference; see Fig. 8. The related methods are as follows.
Figure 8: An illustration of the taxonomy of (a) regularization-based approach and (b) its respective methods; adapted from studies in [16].
  1. 1.
    Weight Regularization: It includes techniques such as Elastic Weight Consolidation (EWC) [67], Synaptic Intelligenc (SI) [63], Memory Aware Synapses (MAS) [143], Riemannian Walk (RWalk) [144]. Meanwhile, other techniques are categorized into Expansion-Renormalization [95, 145, 146, 87], Quadratic Penalty Refinement [147, 148, 149], and Online Variational Inference [150, 151, 152, 153, 154, 93, 155].
  2. 2.
    Function Regularization: It regulates the variations in models’ function $f_{\theta}$ over time, and applies constraints directly to the function. A generalized form of function regularization can be expressed as:
    $\theta_{t}=\arg\min_{\theta}\left(\frac{1}{N_{t}}\sum_{i=1}^{N_{t}}\mathcal{L}_{t}(f_{\theta}(x_{i}),y_{i})+\lambda\mathcal{R}_{f}(f_{\theta},\mathcal{D}_{1:t-1})\right)$ (3)
    where, $\theta_{t}$ are the model parameters after learning task $t$, $\mathcal{L}_{t}$ is the loss function for task $t$, $N_{t}$ is the number of samples in task $t$, $\mathcal{R}_{f}(f_{\theta},\mathcal{D}_{1:t-1})$ is the function regularization term that penalizes changes in the models’ function based on its performance on previous tasks $\mathcal{D}_{1:t-1}$ and $\lambda$ is a hyperparameter controlling the strength of the regularization.
  3. 3.
    Knowledge Distillation (KD): It utilizes the previously-trained model as the teacher and the currently-trained model as the student [16]. It includes techniques studied in [156, 157, 158, 159, 160].
  4. 4.
    Besides the above prominent methods, several alternatives have also been proposed: Sequential Bayesian Inference over Function Space [16, 161, 162, 163] and Conditional Generation [164, 165, 166].
II-E3 Rehearsal/Replay-based Approach
Figure 9: An illustration of the taxonomy of (a) replay-based approach and (b) its respective methods; adapted from studies in [16].
This approach is characterized by approximating and recovering old data distributions [16]; see Fig. 9. The related methods are discussed in the following.
  1. 1.
    Experience Replay (ER): It stores a few old training samples in a small memory buffer. It includes several strategies such as Sample Selection (SS) [85, 167, 86, 157], Gradient/Optimization-based Techniques [16, 103, 168, 169, 170, 171, 172], Storage Efficiency [173, 174, 174, 65, 171, 175, 176], Integration of ER with KD [157, 177], Mitigation of Data Imbalance [178, 179, 180, 180, 165, 181, 112, 182, 183, 184, 185], Learning Plasticity Enhancement [186, 187], and Overfitting Alleviation [188, 189, 190, 191].
  2. 2.
    Generative Replay: It uses an additional generative model such as Generative Adversarial Networks (GANs) [192] or Variational Autoencoders (VAEs) [193] to replay the previously-learned data [21]. There are several GANs-based techniques as studied in [194, 164, 64, 195, 196]. However, they usually suffer from label inconsistency [64, 197]. Toward this, other techniques employ autoencoder-based strategies [198, 197, 21].
  3. 3.
    Feature Replay: It includes techniques studied in [22, 23, 199, 200, 201, 202, 203, 204].
II-E4 Optimization-based Approach
This approach is characterized by explicitly designing and manipulating the optimization programs [16]; see Fig. 10. It includes the following methods.
Figure 10: An illustration of the taxonomy of (a) optimization-based approach and (b) its respective methods; adapted from studies in [16].
  1. 1.
    Gradient Projection: It includes techniques studied in [205, 206, 207, 208, 209, 210, 211, 212, 213, 214] and the ones with Replay-based Strategies [86, 215, 216, 85].
  2. 2.
    Loss Landscape: The related techniques are studied in [217, 218, 219].
  3. 3.
    Meta-Learning: It is also known as learning-to-learn for CL, which attempts to obtain a data-driven inductive bias for various scenarios, rather than designing it manually [15]. It includes techniques studied in [220, 221, 222, 85, 223, 224, 225, 226, 227, 228, 229].
II-E5 Architecture-based Approach
This approach is characterized by constructing task-specific parameters, that can explicitly resolve the inter-task interference caused by incremental task learning with shared parameters [16]; see Fig. 11. It encompasses the following methods.
Figure 11: An illustration of the taxonomy of (a) architecture-based approach and (b) its respective methods; adapted from studies in [16].
  1. 1.
    Parameter Allocation: It isolates parameter subspace, which is dedicated to each task throughout the model, where the architecture can be fixed or dynamic in size. It includes strategies like Fixed Architecture [66, 94, 155, 230, 231, 232, 233, 93] and Dynamic Architecture [234, 235].
  2. 2.
    Modular Network: It leverages parallel sub-networks to learn new tasks one-by-one, without pre-defined task-sharing or task-specific components. The related techniques are studied in [236, 237, 90, 238, 239, 240, 241, 242, 243].
  3. 3.
    Model Decomposition: It separates a model into the task-sharing and task-specific components [16] (Fig. 12). The relate techniques are studied in [244, 245, 246, 247, 248, 249, 250, 251].
    Figure 12: An illustration of the model decomposition method exhibiting two types; corresponding to parameters (i.e., low-rank factorization) and representations (i.e., masks of intermediate features); adapted from studies in [16].
All these existing methods for fulfilling the desiderata of CL are also summarized and presented in Table III.
TABLE III: Summary of CL methods based on Representation-, Regularization-, Rehearsal/Replay- and Architecture-based approaches that potentially fulfill “+” or have fulfilled “✓” the desiderata of CL.
Desiderata of CL
Approach Methods Work,(Year)
Scalability No/Minimal Task Positive Positive Controlled Fast
Old Data Use Agnostic FWT BWT Forgetting Adaptation
[69],(2023)
Pre-training for [252],(2023)
Downstream [73],(2024)
Tasks [253],(2024)
MAML [98],(2017)
Self-Supervised [70],(2023)
Representation Learning
[254],(2019)
[79],(2022)
Continual TADIL [81],(2023)
Pre-training [255],(2022)
AFEC [87],(2021)
CFN [96],(2020)
GPM [208],(2021)
[68],(2022)
[256],(2022)
[74],(2018)
Weight [257],(2023)
Regularization OGD [206],(2020)
Regularization TRGP [210],(2022)
OWM [205],(2019)
ASP [97],(2017)
UCL [93], (2019)
Function HCL [77],(2021)
Regularization SGP [258],(2023)
Distillation [259],(2023)
[82],(2022)
Experience [260],(2019) +
Replay Replay MER [85],(2019) +
GEM [86],(2017)
Generative TAME [78],(2024)
CN-DPM [76],(2020)
Model DSSAE [80],(2023)
Architecture Decomposition P&C [95],(2018)
CLNP [94],(2019)
Allocation AFAF [91],(2022)
II-F CL Settings
II-F1 Supervised Learning
It is an ML setting that trains models with labeled datasets. Therefore, in the Supervised CL setting, the goal is to train a model on a sequential data stream. Its challenge is to efficiently perform the training, since the supervised-based learning typically requires huge memory and energy requirements. Moreover, obtaining a large labeled dataset can be time consuming and expensive.
II-F2 Unsupervised Learning
It is an ML setting that trains models with unlabeled data (e.g., identifying patterns, structures, or relationships) [261]. Therefore, in the Unsupervised CL setting, the goal is to develop a model that can discover similarity in samples that arrive sequentially [261]. Its challenge is to extract meaningful features without explicit labels. This setting is particularly useful for tasks where data labeling is impractical or expensive.
II-F3 Reinforcement Learning (RL)
It is an ML setting that trains models to make sequential decisions based on the feedback (i.e., rewards or penalties) [262, 263]. Through trials and errors, the model learns to associate actions with outcomes, aiming to discover the optimal strategy or policy to achieve its objectives over time. Its challenge is to efficiently perform the training, as RL typically requires high memory and energy requirements. Moreover, the nature of trials and errors can also be time-consuming and expensive.
II-G Evaluation Metrics
To evaluate the performance of an efficient CL system, metrics are essential. To assess the overall performance of a CL algorithm, average accuracy (AA) [264, 86] and average incremental accuracy (AIA) [157, 178] are often used. Model size is used to measure the memory size for storing the model parameters [265], while sample storage size (SSS) is used to measure the storage size for storing data samples. To evaluate the computational resources required for the training and inference phases, metrics such as the number of floating-point operations-per-second (FLOPS) and time complexity are often used. Meanwhile, power/energy efficiency measures the efficiency gains of a CL system during its operational life-time.
III Neuromorphic Computing
III-A Overview of SNNs
Spiking Neural Networks (SNNs) mimic the brains’ functionality through the utilization of spikes for transmitting information [37, 266, 267]. Therefore, SNNs encode information into discrete spike trains. Popular encoding techniques include rate coding [268] (e.g., spike count, spike density, population activity [269]), and temporal coding (e.g., burst [270], time-to-first spike (TTFS) [271], phase [272], and rank-order [273]). Each spiking neuron processes the input spikes, and its behavior (i.e., neuronal dynamics) depends on the neuron model, such as Hodgkin–Huxley (HH) [274], Leaky Integrate and Fire (LIF) [275], Izhikevich Model [276], and Adaptive Exponential Integrate-and-Fire (AdExIF) [277]. Neuron model selection typically considers the expected neuronal dynamics and computational complexity [278, 279]. The output spikes are generated only when neurons’ membrane potential reaches the threshold, and transmitted through synapses, enabling ultra-low power/energy consumption [52, 280, 281, 282, 283].
SNN training can be performed using bio-plausible or analytical learning rules. The bio-plausible ones are typically characterized by local learning mechanism, like Hebbian [284], Spike-Timing-Dependent Plasticity (STDP) [285], Spike-Driven Synaptic Plasticity (SDSP) [286, 287], and Reward-Modulated STDP (R-STDP) [288, 289]. Meanwhile, the analytical ones encompass DNN-to-SNN conversion [290], Surrogate Gradient Learning [291] (e.g., Backpropagation Through Time (BPTT) [291, 292, 293], Spatio-Temporal Backpropagation (STBP) [294], and Deep Continuous Local Learning (DECOLLE) [295]), Spike-Triggered Local Representation Alignment (ST-LRA)  [296], and Bayesian Learning [297].
SNN processing demands a suitable computing hardware to maximize its potentials in accuracy, latency, and power/energy efficiency. Conventional von-Neumann architecture-based hardware platforms (e.g., CPU and GPUs) have been widely used to perform Python-based SNN processing using generic arithmetic units  [298, 299, 300, 301], thereby leading to sub-optimal efficiency gains. To address this, CMOS-based neuromorphic hardware accelerators facilitate efficient spike transmission and computation [302], allowing their implementation in the Field-Programmable Gate Array (FPGA) or Application-Specific Integrated Circuit (ASIC). Popular accelerators include Neurogrid [303], ROLLS [304], TrueNorth [305], and Loihi [306]. Beyond CMOS-based technologies, the processing-in-memory (PIM) or compute-in-memory (CIM) paradigm has been explored to further reduce latency and energy in data transfer between memory and compute units [307, 308] by leveraging non-volatile memory (NVM) technologies, such as Resistive Random Access Memory (RRAM), Magnetic RAM (MRAM), and Phase Change Memory (PCM) [309].
III-B Neuromorphic Continual Learning (NCL)
NCL is an advancement of bio-plausible SNNs for facilitating continual adaptation in dynamic environments. Here, we discuss the state-of-the-art NCL methods, including their key ideas and limitations, whose details are summarized in Table IV.
III-B1 Enhancements on Unsupervised STDP Learning
While conventional STDP enables efficient online unsupervised learning, employing STDP alone may still suffer from CF [41, 42]. To address these challenges, proposed techniques have been proposed in the literature, as discussed in the following.
Weight Decay: It prevents overfitting in NNs by penalizing large weight values by adding a regularization term, which encourages the model to keep the weights small. For instance, Adaptive Synaptic Plasticity (ASP) [97], SpikeDyn [41] and lpSpikeCon [42] leak the weights to gradually remove old and insignificant information and ensure the weights do not grow too large.
Adaptive Threshold Potential: It refers to the dynamic adjustment of the neurons’ threshold based on its spiking activity. It allows neurons to adapt their sensitivity to incoming spikes, making the network more flexible/robust in processing information. A neuron typically fires when its membrane potential exceeds a pre-defined threshold. However, in an adaptive threshold model (as employed by SpikeDyn [41] and lpSpikeCon [42]), the threshold changes based on the neurons’ recent activity, thus helping the neuron provide the following features.
  1. 1.
    Regulating the firing rates to prevent excessive firing and maintain a stable firing rate over time to avoid overactive neurons (homeostasis[268, 266].
  2. 2.
    Leveraging the temporal information to improve the neurons’ capability to respond to varying input patterns.
  3. 3.
    Improving the signal-to-noise ratio (SNR), by separating the significant input patterns from the noise.
Adaptive Learning Rate: It dynamically adjusts the learning rate in training to improve convergence. The learning rate controls how much the models’ weights are updated. ASP [97], SpikeDyn [41] and lpSpikeCon [42] employ adaptive learning rates to determine the potentiation and depression factors in the STDP-based learning based on the spiking activities. By prioritizing the adjustment of highly active synapses, it enhances learning efficiency and reduces unnecessary energy expenditure on less active connections. This selective learning process mimics biological synaptic behavior.
These techniques are often employed together in NCL methods as discussed below.
Adaptive Synaptic Plasticity (ASP) [97]: It integrates weight decay with the STDP-based weight updates to balance forgetting and learning, while leveraging time-dependent learning rate. ASP uses a two-phase weight update process (i.e., recovery and decay), allowing weight updates based on spiking activities and gradually leak toward a baseline value; see Fig. 13. Despite its strengths, ASP requires larger quantities of input samples from earlier distributions than later ones, hence requiring the knowledge of task changes [96], making it unsuitable for OCL scenario.
Figure 13: Weight update process of ASP; adapted from studies in [97]. More frequent input spikes represent common features between old and new input patterns, and will experience a greater weight update compared to less frequent spikes, that correspond to unique features of a specific input.
Controlled Forgetting Network (CFN) [96]: It is an SNN architecture that exploits dopaminergic neurons to modulate the synaptic plasticity. Its idea is to temporarily make the weights of some neurons more plastic (easier to change) and keep the weights of other neurons. The modulation is triggered by dopaminergic neurons, which fire when there are no or only a small number of incoming inhibitory spikes (indicating that new information is encountered), and then stimulate a temporary boost of learning rate for other neurons (Fig. 14). Despite its benefits, CFN requires additional components (i.e., dopaminergic neurons and their connections) and still considers a simple MNIST dataset with conventional CL scenario, hence making it challenging for deployments with more complex datasets under OCL scenario.
Figure 14: Single-layer CFN architecture; adapted from studies in [96]. The dopaminergic neuron activates when other neurons in its layer are inactive, indicating new information in the input. This activation temporarily enhances plasticity in other neurons. Dopaminergic signals are weighted to provide targeted stimulation.
SpikeDyn [41]: It focuses on enhancing STDP for enabling unsupervised CL while minimizing energy consumption in both training and inference phases (Fig. 15). Its key ideas include: reduction of neuronal operations by substituting the inhibitory neurons with direct lateral inhibitions to reduce memory and energy requirements; enhancing unsupervised CL algorithm by employing adaptive learning rates, weight decay, adaptive threshold potential, and reduction of spurious weight updates; and SNN model size search by leveraging analytical models to estimate the memory and energy requirements, and selecting a Pareto-optimal model that meets the resource constraints. SpikeDyn demonstrated that optimization techniques can be exploited to minimize memory and energy requirements of CL systems, but they should be supported with an enhanced unsupervised CL algorithm to maintain the performance.
Figure 15: SpikeDyn framework for adaptive and energy-efficient unsupervised CL for SNNs; adapted from studies in [41].
lpSpikeCon [42]: It extends the studies in SpikeDyn [41], by enabling STDP-based unsupervised CL under low-precision settings for autonomous embedded systems (e.g., robots). lpSpikeCon employs weight quantization and compensates the loss of information by identifying and adjusting SNN parameters that significantly impact accuracy (Fig. 16). This study found that the key SNN parameters to adjust include the adaptive threshold potential and weight decay rate. This study demonstrated that low-precision settings can be exploited for substantially reducing memory and energy requirements, and adjustments of other parameters are crucial to maintain the performance.
Figure 16: Key steps of lpSpikeCon methodology for low-precision unsupervised CL for SNNs; adapted from studies in [42].
Despite their benefits, SpikeDyn and lpSpikeCon still consider a simple MNIST dataset and conventional CL scenario, thus requiring further studies for OCL scenario.
III-B2 Predictive Coding
This method is used by the Spiking Neural Coding Network (SpNCN) [296] to predict incoming data and then correct the prediction based on the actual inputs [310, 311]; see Fig. 17. This iterative process of “guess-and-check” allows the network to adjust its weights continually and learn from data streams without repeated exposure to same data. Key ideas of SpNCN are as follows.
Figure 17: A 2-layer SpNCN with error units (blue diamonds) that measure the difference between predictions ($z^{0}_{\mu}$, $z^{1}_{\mu}$) and target signals ($z^{0}(t),z^{1}(t)$). Variables $s^{0}(t),s^{1}(t)$ are the binary spike outputs of neuron groups at time $t$. Black dash-dotted arrows show the repeated transmission of the last known values. Mismatch signals, shown by green dashed arcs, adjust the spiking neuron’s action potentials through synapses. Solid black lines show predictive synapses, and black dotted lines show direct information transfer; adapted from studies in [296].
  • Prediction of neuron activity and error correction using local synaptic updates.
  • Weight adaptation using a coordinated ST-LRA, which adjusts weights based on the mismatch between predictions and actual activity. It is also combined with STDP as regularizer.
  • A memory module and task/context-modulated lateral inhibition to enhance memory retention across tasks.
This study showed that combination of learning rules is potential for enabling CL. However, this complex combination of multiple learning rules make it challenging for deployments under OCL scenario.
III-B3 Active Dendrites
Studies in [312] proposed an SNN model leveraging active dendrites [313] to facilitate task-specific sub-network formation (Fig. 18). The spike time $t_{j}$ of a neuron is modulated by a function of the selected dendritic segment $u_{jn}$ for the current task. The dendritic activation function $f(u)$ modulates the spike time dynamically, allowing the model to adapt its behavior based on the task context. It also leverages TTFS encoding to introduce a gating mechanism, that enables selection of sub-networks for various tasks. However, this method requires labeled data or supervisory signals, hence making it challenging for deployments under the OCL scenario.
Figure 18: Neuron model and network architecture with active dendrites; adapted from studies in [312]. (a) Linear integration of synaptic strength $W_{ij}$ following a pre-synaptic spike at $t_{i}$; Top figure shows the modulation of spike timing by dendritic processes. (b) Illustration of a pyramidal neuron. (c) Selection of different sub-networks for various tasks based on dendritic segment activity. (d) Proposed neuron model and dendritic activation function.
III-B4 Bayesian Continual Learning
This method represents weights with parameters that quantify the epistemic uncertainty based on prior knowledge and observed data, and employs Bayesian methods for handling uncertainty over time by determining which knowledge to retain and which to forget [297] (Fig. 19). For real-valued synapses, it uses a Gaussian variational distribution to adjust the values; while for binary synapses, it uses a Bernoulli variational distribution with Gumbel-softmax [314]. Although, this method provides better-calibrated decisions and better detection compared to conventional frequentist approaches [315], the uncertainty quantification incurs high computational complexity, thereby making it challenging for deployments in OCL scenario.
Figure 19: In the Bayesian continual learning, the system is sequentially presented with similar yet distinct tasks; adapted from studies in [297].
III-B5 Architecture-based Approach
Dynamic Structure Development of Spiking Neural Networks (DSD-SNN) enhances the SNN structure by growing new neurons for new tasks and pruning redundant neurons [316] (Fig. 20). It employs a deep SNN architecture comprising of multiple convolutional and fully-connected (Fc) layers, which is equipped with random growth, adaptive pruning, and freezing of neuron mechanisms. This study advanced multi-task learning while enhancing memory capacity and efficiency. However, the development of dynamic structure adds complexity to SNN implementation, which is challenging for OCL scenario.
Figure 20: Deep SNN architecture with dynamic structure development; adapted from studies in [316].
Self-Organized Regulation SNN (SOR-SNN) employs a pathway search module to adaptively activate task-specific sparse neural pathways based on fundamental weights $W_{t}$ [317] (Fig. 21). Each synapse has two states (i.e., active and inactive), and is determined by learnable synaptic selection parameters. The model decides whether to activate or inhibit each weight by comparing the learnable parameter $A_{s}$ with the threshold $\tilde{A}_{s}$, where activation is preferred if $A_{s}>\tilde{A}_{s}$. This study showed that SOR-SNN enables adaptive reorganization, BWT, and self-repair capabilities. Despite its benefits, this method requires complex dynamic connectivity capabilities, hence making it challenging for deployments under OCL scenario.
Figure 21: The process of SOR-SNN model; adapted from the studies in [317]. Each SNN block includes a self-organizing regulation network which selectively activates task-specific sparse pathways. For instance, the purple connections represent the pathway for Task 1. The numerous combinations of connections enable the limited SNN to incrementally learn a larger number of tasks.
III-B6 Rehearsal/Replay-based Approach
A memory replay approach using Experience Replay (ER) method has been developed, where a convolutional SNN is trained to learn in Class-IL and TA scenarios [318]. For resource-constrained devices, the memory-efficient Latent Replay (LR)-based method has been developed, which stores compressed data representations [292]. The Latent Replay training involves a pre-training phase, where SNN is divided into two parts i.e., frozen and learning layers (Fig. 22). The network stores latent replays for memory, and only trains the learning layers on new data. Here, the main challenges are related to the efficient store-load mechanisms for replay data under OCL scenario.
Figure 22: SNN with Recurrent Neurons and Latent Replays; adapted from studies in [292]. (a) Recurrent neuron model. (b) Structure of a generic fully-connected model for CL. (c) An example of the data fed to the Latent Replay stage.
III-B7 Regularization-based Approach
Figure 23: The effective potential force field ($U^{*}$) created by the R-STDP mechanism, which prevents the Brownian motion (represented by the dashed line) from leaving the optimal weight domain $D$; adapted from studies in [289].
This approach employs regularization terms to balance the old and new tasks, such as Noise Regularization [319], Freezing Large Weights, and Stochastic Langevin Dynamics [289]. For instance, Langevin Dynamics exploit the fact that each weight $w_{i}$ can vary without impacting the accuracy in the domain $D_{i}$ if constrained to a specific space when learning new tasks; see Fig. 23. This study highlighted that Langevin dynamics can exhibit very similar performance as the replay-based methods while being less memory-intensive. However, its compute requirements are relatively high, thus requiring further studies for OCL scenario.
III-B8 Hebbian Learning
Hebbian Learning-based Orthogonal Projection (HLOP) method leverages lateral connections and Hebbian learning to achieve CL [320]. It employs orthogonal gradient projection to modify pre-synaptic activity traces (Fig. 24) to ensure that weight updates for new tasks do not interfere with the weights associated with the old tasks. Hebbian learning is useful to extract the principal subspace of neural activities, and updates synaptic weights based on the correlation between pre- and post-synaptic neuron activities, while the activity traces are updated using lateral signals.
Figure 24: An illustration of the HLOP; adapted from studies in [320]. (a) Overview of the method. (b) Skew-symmetric connection weights in lateral circuits. (c) Construction of new subspaces for new tasks with Hebbian learning in lateral circuits. (d) Recurrent lateral connections based orthogonal projection.
III-C Efficiency Enhancement Methods for NCL
To reduce the memory footprint and energy consumption of SNNs with CL capabilities (NCL), several optimization methods have been proposed and discussed below.
III-C1 Reduction of SNN Operations
Real-time resource-constrained applications require fast and efficient processing SNNs. Their operations can be reduced through several network optimizations by:
  • Leveraging sparse neurons and sparse synapses to reduce the number of neuron operations and connections/weights, respectively.
  • Utilizing temporal encoding uses timing of spikes to encode information in a way that can reduce computational complexity compared to rate encoding.
  • Employing simple neuron models like LIF instead of complex models to reduce the computational complexity required per neuron.
SpikeDyn [41] replaced the inhibitory neurons with the direct lateral inhibitions in the SNN model used in ASP [97], which consisted of input, excitatory, and inhibitory layers (Fig. 25), thereby eliminating the operations in the inhibitory layer.
Figure 25: Replacement of the inhibitory neurons with direct lateral inhibitions; adapted from studies in [41].
III-C2 Weight Quantization
This method enhances SNNs efficiency by representing network weights with lower precision instead of using full-precision floating-point numbers [321, 281, 282, 283, 322, 323, 324]. Weights can be quantized to a smaller number of selected discrete levels such as 8-bit, 4-bit, or even binary, using uniform or non-uniform quantization [325]. For instance, lpSpikeCon [42] reduced SNN weights using uniform quantization with truncation [326, 325]. Similarly, to reduce the required on-chip memory, the synaptic weights and dendritic delays are quantized in [313]. Thus, keeping the user-defined number of bits and removing the remaining bits from the fractional part, reduces the storage size, speeds up the training and inference processes, and enables more efficient hardware implementations [266].
TABLE IV: Summary of the state-of-the-art works for energy-efficient CL employing bio-plausible SNNs (i.e., NCL).
Work, CL Learning Neuron Neural Learning Learning Optimization Dataset Tools
Year Scenario Setting Model Coding Rule Rate Technique
ASP [97], Class-IL Unsupervised LIF Rate STDP Adaptive Weight Decay MNIST Brian
(2017) [327]
SpikeDyn [41], Class-IL Unsupervised LIF Rate STDP Adaptive Weight Decay MNIST Python
lpSpikeCon [42], Class-IL Unsupervised LIF Rate STDP Adaptive Weight MNIST Python
(2022) Quantization [299]
CFN [96], Class-IL Unsupervised LIF Rate STDP Adaptive - MNIST
SpNCN [296], Task-IL Supervised LIF Rate ST-LRA, - - MNIST [328]
(2023) ST-LRA+ FMNIST [329]
STDP Not-MNIST [330]
[312],(2024) Task-IL Supervised Enhanced TTFS Fixed Weight and Delay Split-MNIST FPGA,
model of Quantization Python
[331]
[297],(2022) Class-IL Supervised SRM Rate Bayesian - - Split-MNIST, Intel’s
Learning MNIST-DVS Lava
[332] [333]
DSD-SNN Task-IL, Supervised LIF - Back- - - Permuted MNIST BrainCog
[316],(2023) Class-IL Propagation Not-MNIST [330], [334]
Split-CIFAR100
SOR-SNN Class-IL Supervised LIF - Back- Adaptive - CIFAR100 [335], -
[317],(2023) Propagation ImageNet [336]
[318],(2023) Class-IL, Supervised LIF BPTT Fixed MNIST PyTorch
TA Fixed [337]
[292],(2023) Sample-IL, LIF - BPTT - Time SHD
Class-IL Supervised Compression [338]
[289],(2022) Task-IL Unsupervised IF Rank- STDP, $a^{+}<0.15$, - MNIST, SpykeTorch
Order R-STDP $a^{-}>-0.1125$ EMNIST [339] [340]
HLOP-SNN Task-IL, Supervised LIF Rate Hebbian Fixed - PMNIST [328], Python
[320],(2024) DIL CIFAR100
III-D Summary of NCL
The state-of-the-art methods have performed initial studies for enabling NCL (Table IV). We observe that methods with bio-plausible learning rules (e.g., STDP) facilitate unsupervised learning, which is suitable for OCL scenario. These works mainly employ fully-connected SNN architectures (see Table V), consisting of input layer and a pair of excitatory-inhibitory layer. As the input data grows in complexity, fully-connected SNNs struggle to capture the important hierarchical features without increasing the network size. Larger architectures often lead to better accuracy but require larger storage sizes, which may exceed the resource constraints. This limits the deployable networks on practical hardware platforms, and make their performances limited to simple datasets (e.g., MNIST) with simple CL scenarios (e.g., Split-MNIST). Meanwhile, methods that employ supervised-based learning rules (e.g., surrogate gradient-based backpropagation) can achieve acceptable performance on complex datasets (e.g., ImageNet). However, they are not suitable for OCL scenario, as they require costly labeled data and power-hungry training. Furthermore, these works mainly focus on energy-efficiency, and have not optimized resource utilization. Therefore, alternative methods are required especially for enabling energy-efficient SNNs with OCL capabilities, which are beneficial for many real-world application use-cases.
TABLE V: Comparative analysis of the state-of-the-art in terms of network architecture, model size, performance, memory footprint, and power/energy savings. Note, FC: fully-connected layer, and CONV: convolutional layer.
Work SNN Model Performance Memory Power/Energy
Architecture Size Footprint Consumption
ASP [97] 2 FC layers 200 excitatory & Achieved avg. accuracy of 75.8% (exp. decay) n/a n/a
(1 input layer & 200 inhibitory and 74.5% (linear decay), outperforming STDP
1 excit.-inhib. layer) neurons by $\sim$5.7%, which had 70.1% accuracy
SpikeDyn 2 FC layers (1 input 200 and 400 Improved accuracy by 21% for the last learned $<$3000 KB Avg. reduction of 51%
[41] & 1 excit. layers) excit. neurons task and 8% for previous tasks than ASP (training) and 37%
(inference) than ASP
lpSpikeCon 2 FC layers (1 input 200 and 400 No accuracy loss in inference compared to 8x reduction n/a
[42] & 1 excit. layers) excit. neurons the 32-bit weights baseline (SpikeDyn) from SpikeDyn
CFN Multi-FC layers 400-6400 Achieved 95.24% avg. accuracy across MNIST n/a n/a
 [96] (1 excit., 1 inhib., & excit. neurons digits in Class-IL for 6400-neuron CFN
1 dopaminegic layers)
SpNCN 4 FC layers 3000 neurons- Normalized accuracy of $0.9653$ (MNIST), n/a n/a
 [296] per-layer $0.9120$ (Not-MNIST), and $0.9995$ (FMNIST)
[312] 4 FC layers 784-403-403-2 Achieved avg. inference time of 37.3ms 35.3% flip-flops, n/a
784-400-400-2 & test accuracy of 80.0% on FPGA 29.3% BRAMs
[297] Multi-FC layers 2048, 4096, 4096 Achieved avg. test accuracy of 85.44±0.16% 5× reduction n/a
2048, 1024 neurons for Split-MNIST and 74% for MNIST-DVS
per-layer
DSD-SNN Multi-CONV & 100 & 500 Achieved 97.30%±0.09% accuracy (34.38% n/a n/a
[316] 1 FC layers neurons compression rate) for MNIST in Task-IL,
96.94%±0.05% accuracy for N-MNIST,
60.47% (10 steps) for CIFAR100 in Class-IL
SOR-SNN ResNet18 96 hidden Achieved 86.65%±0.20 accuracy, & improved n/a 0.32M parameters
[317] layer neurons accuracy by 2.20% from DSD-SNN (CIFAR100) (9.35% of DSD-SNN)
[318] 2 CONV layers n/a Achieved 51% avg. accuracy (40 batches-per-task) n/a n/a
in Class-IL & 36% accuracy with memory replay
(100 batches-per-task) in Task-Free scenario
[292] 4 Recurrent-FC 200-100-50-20 Achieved Top-1 accuracy of 92.5% and 92% in Two orders of n/a
layers neurons Sample and Class-IL scenarios, respectively magnitude reduction
and 78.4% in multi Class-IL scenario from naive rehearsal
(max. 4% acc. drop)
[289] 3 CONV layers 30-250-200 Achieved 92.0±0.1% accuracy for digits in n/a n/a
neurons joint training, & 79.7±0.5% accuracy for letters
HLOP-SNN FC, ResNet-18, & 784-800-10 Achieved 95.15% accuracy (PMNIST), n/a n/a
[320] 3 CONV layer neurons for 78.58% (CIFAR-100), 63.40% (miniImageNet),
architectures FC SNN & 88.65% (5-Datasets)
IV Real-World Application Use-Cases
In this section, we discuss the real-world application use-cases that benefit from bio-plausible SNNs with OCL capabilities. Additionally, we also highlight the current state-of-the-art for these application use-cases.
IV-A Adaptive Robots
IV-A1 Object Recognition for Humanoid Robots
Adaptive Robots need to identify, classify, and locate objects within an image/video. For this, Neural Engineering Framework (NEF) is explored for developing Neural State Machine (NSM) framework leveraging SNNs, managing the allocation of new neurons (Fig. 26), where data may be collected on-the-fly without any labels [341]. However, it relies on supervised settings which require costly labeled data, power-hungry training, and long training time to update its knowledge, thus it is not suitable for OCL scenario.
Figure 26: The NSM framework: Subsampling layer (SS), S1 and C1 perform the visual feature extraction, S2 is the plastic layer (dark blue neurons) that stores visual patterns, Error (Err) and label neurons (Lb) control inference decision of output neurons (Out); adapted from studies in [341].
IV-A2 Inverse Kinematics and PID Control
Recent study explored Neural Engineering Framework (NEF) for developing neuromorphic algorithms for Inverse Kinematics (IK) and Proportional-Integral-Derivative (PID) control systems, supporting online learning of control signals, using Prescribed Error Sensitivity (PES) [342]. Fig. 27(a) shows the model schematic for neuromorphic IK with online learning, whose target is to efficiently compute the joint angles required to place the robots’ end-effector at the desired position. Meanwhile, the PID controller aims to ensure that the robots’ actuators accurately follow a desired trajectory. Hence, it needs to continuously adjust the actuators to minimize the error in position, ensuring accurate motion control. To achieve this, the PID controller has been implemented using spiking neurons to represent the error signals and components; see Fig. 27(b). However, it relies on supervised settings which require costly labeled data, power-hungry training, and long training time to update its knowledge and thus is not suitable for the OCL scenario.
Figure 27: (a) A model for neuromorphic IK with online learning, and (b) Neuromorphic PID controller in action; adapted from studies in [342].
IV-B Autonomous Vehicles and Mobile Agents
CL in autonomous vehicles and mobile agents is crucial to enable continuous adaptation to the changes in environments (e.g., road and weather) and personalize the preferences. Hence, they are typically equipped with sensors and cameras for continuous data collection during operation. Moreover, they usually rely on limited battery, and often face unlabeled data and constrained resources. Therefore, their use-cases can benefit from SNNs with OCL capabilities.
IV-B1 Cars Detection
Several works have been developed for low-power autonomous vehicles using SNNs, including studies of cars detection in CarSNN [343], SNN4Agents [324], FastSpiker [344] and [345, 346]. CarSNN and SNN4Agents used an attention window mechanism (i.e., focusing on regions with the highest event density) to process event-based input samples from the N-CARS dataset [347]; see Fig. 28. However, the state-of-the-art [343, 324] require a long training time [344]. Therefore, FastSpiker proposed a methodology to accelerate SNN training while maintaining accuracy through learning rate enhancements [344]. However, these works considered supervised settings which require costly labeled data, power-hungry training, and long training time to update its knowledge, hence they are not suitable for OCL scenario.
Figure 28: Setup and tool-flow of CarSNN; adapted from studies in [343].
IV-B2 Road Lane Detection
Other SNN works for low-power autonomous vehicles include the road line detection. For instance, LaneSNN [348] addressed the imbalance between lane and background classes by employing a novel loss function combining Weighted Binary Cross Entropy (WCE) and Mean Squared Error (MSE); see Fig. 29. Evaluations demonstrated comparable performance to state-of-the-art methods like LaneNet and RefineNet but with significantly lower power consumption. However, it still considered supervised settings which require costly labeled data, power-hungry training, and long training time to update its knowledge, hence it is not suitable for OCL scenario.
Figure 29: Setup and tool-flow of LaneSNNs; adapted from studies in [348].
IV-B3 Simultaneous Localization and Mapping (SLAM)
Figure 30: DVS-Radar Fusion for SNN-STDP based CL SLAM system; adapted from studies in [349].
SLAM is important for an autonomous vehicle/agent for constructing a map while simultaneously keeping track of its location in an unknown environment. The DVS-Radar Fusion SLAM [349] is the first SNN-based method for enabling CL-based SLAM using drones (Fig. 30), integrating DVS and a Frequency Modulated Continuous Wave (FMCW) radar to encode sensory data on-the-fly into feature descriptors. To provide drones’ velocity information and aid in obstacle detection, a radar-gyroscope odometry method is used for accurate navigation and mapping. SNN output is utilized as feature descriptors for loop closure detection that are fed to a RatSLAM back-end. This process helps in identifying when the drone revisits a previously mapped area. The obstacle modelling is achieved using radar data by detecting objects and integrating this information into its SLAM framework. This work showed the potential of implementing OCL on SNN-based mobile agents.
IV-B4 People Detection
A CL-based people detection framework has been developed in [350]; see Fig. 31. It employs attention maps to enable network adaptation on walking-people detection in dynamic environments, thus avoiding collision. To create attention maps from DVS data, the network is integrated with a readout mechanism. It also generalizes the unsupervised STDP from [351] to a semi-supervised learning case with teacher signals (TSs). To control overfitting, an anti-Hebbian rule (i.e., negative STDP) is applied when person is absent, while positive STDP is used when a person is detected. This work further showed the potential of implementing OCL on SNN-based systems.
Figure 31: Attention maps based CL system for people detection; adapted from studies in [350]. (a) The DVS produces a stream of spikes over time and space. (b) A spike is emitted when the change in light intensity $L_{ij}$ at pixel $(i,j)$ exceeds a certain threshold. The spike is positive if the change $\Delta L_{ij}>0$ and negative otherwise. (c) The SNN-STDP followed by a readout of neural activity is used to investigate the continuous development of attention-based perception.
A summary of real-world application use-cases that will benefit from SNNs with OCL, is presented in Table VI.
TABLE VI: Summary of real-world application use-cases that will benefit from SNNs with OCL capabilities, and their current state-of-the-art works.
Applications Work CL Dataset CL Learning Neuron Hardware
(Year) Setting Approach Rule Model
Object [341] Supervised Custom Representation- Modified LIF Loihi
Recognition (2022) DVS dataset based Hebbian
Adaptive
Robots Custom
IK and PID [342] Supervised dataset from Representation- PES LIF Loihi
Control (2021) robotic arms’ based
operations
Cars [343] (2021) STBP LIF Loihi
Detection [345] (2022) Supervised N-CARS - Surrogate Gradient LIF GPU
[324] (2024) STBP LIF GPU
Road Lane [348] Supervised DET - STBP LIF Loihi
Detection (2022) [352]
Autonomous
Vehicles/
Agents SLAM [349] Unsupervised Real-time data Representation- STDP LIF n/a
(2023) from sensors based
People [350] Semi-supervised KUL- Representation- STDP LIF ReckOn
Detection (2024) UAVSAFE [353] based [354]
V Open Research Challenges
Due to the huge potentials and benefits of energy-efficient SNNs with OCL capabilities, continuous advancements in the NCL field are expected in the future. Toward this, we identify several open research challenges to be addressed, as follows.
Learning Rule Developments: Leveraging SNN learning rules for enabling OCL capabilities is a major challenge. Such rules may exhibit beneficial characteristics for OCL (e.g., unsupervised local learning), but may also lack some aspects (e.g., context for quick learning convergence). Therefore, developments of appropriate learning rules and their enhancements are critical for achieving OCL capabilities. Furthermore, combining multiple learning rules may also provide promising strategies for solving multiple datasets.
Adapting the Existing CL Methods in SNN Domain: Investigating how existing CL methods in Section II-E can be adapted for SNNs is potential for combining their strengths. For instance, adapting the regularization-based methods and how it impacts the performance of SNNs trained with local learning rules, particularly in dynamic environments could be explored. The optimal memory size for replay-based method in SNNs that balances desiderata effectively, particularly in resource-constrained environments can also be explored.
Balancing CL Desiderata: It is very challenging and often requires trade-offs to satisfy all CL desiderata (i.e., scalability, no/minimal usage of old data, task-agnostic, positive forward and backward transfer, controlled forgetting, and fast adaptation/recovery). For example, improving learning quality for past tasks while minimizing the use of old data is non-trivial. Therefore, it is important to prioritize the desired characteristics based on the applications and balance them during the operational lifetime within acceptable performance.
Generalization Capability: The functionality of SNN-based OCL algorithms should be applicable in broader operational settings. For instance, OCL-based SLAM systems should be able to seamlessly navigate from indoor to outdoor environments. To achieve this, multiple desiderata may need to be satisfied.
Data Acquisition Quality: In OCL scenarios, systems need to learn from unlabelled data from the environments under unsupervised settings [41, 42], which are often noisy. Therefore, input pre-processing and combining multiple high-quality event-based sensors (e.g., DVS camera) with conventional sensors (e.g., LiDAR) are beneficial to enhance the quality of input data.
Hardware Deployments: Deployments of SNN-based OCL algorithms should leverage the inherent low-power and massively-parallel processing strengths of the neuromorphic hardware, thus maximizing the efficiency benefits of neuromorphic computing. For instance, algorithm mapping policies on the given hardware accelerators hold an important role in determining the processing latency and efficiency.
Evaluation and Benchmarks: Standard evaluation protocols and benchmarks are required to ensure fair and comprehensive evaluation metrics for assessing the performance of SNN-based OCL algorithms. These also include different CL scenarios and settings, as well as different metrics like accuracy, model and sample storage size, computation complexity, and power/energy efficiency.
VI Conclusion
This paper provides a comprehensive review of CL foundations, analyzes state-of-the-art CL methods for both DNNs and SNNs (i.e., NCL), highlights the importance of OCL for autonomous embedded systems, then discusses how SNNs can be used for enabling OCL. Specifically, it provides a holistic view on the CL desiderata, scenarios, and settings. Then, it scrutinizes the benefits, and limitations of recent NCL advancements, to highlight the gap between the state-of-the-art and the requirements of practical applications. To address this, this paper identifies open research challenges to enable OCL using SNNs in real-world application use-cases, thereby inspiring future research towards developing energy-efficient CL systems using neuromorphic computing.